# CMRNext: Camera to LiDAR Matching in the Wild for Localization and Extrinsic Calibration

### v2

@@ I Introduction
Localization is an essential task for any autonomous robot, as it is a precursor for subsequent safety-critical tasks such as path planning and navigation. The accuracy of Global Navigation Satellite Systems is not sufficient for most of these tasks, and therefore sensors such as Light Detection And Ranging (LiDAR) and cameras are employed to estimate the pose of the robot. While LiDAR-based localization methods [1, 2] achieve the best performance in terms of accuracy and robustness, the comparatively high price of LiDAR sensors hinders their widespread adoption. On the other hand, cameras are inexpensive and already widely available in consumer vehicles for Advanced Driver Assistance Systems (ADAS) applications. However, the performance of camera-based localization methods [3, 4, 5] is still not comparable to LiDAR-based localization techniques, especially when the lighting and weather conditions during deployment differ from those recorded while mapping.
Consequently, monocular localization in LiDAR maps has gained considerable attention from both academia and industry in recent years, as it combines the high accuracy and robustness of LiDAR-based localization, while only requiring a camera sensor during deployment. The increasing availability of LiDAR maps from mapping companies such as HERE and TomTom has further contributed to the popularity of this approach. Once a LiDAR map is available, either obtained from a mapping company or generated using recent LiDAR Simultaneouos Localization And Mapping (SLAM) algorithms [6, 2, 7], a vehicle can be localized within this map using a single camera. Considering the low cost of cameras, compared to LiDAR sensors, this approach is particularly appealing for large-scale deployment, where the cost of the final system is a key factor. While traditionally this task has been addressed by first generating a 3D representation from the camera image and then matching it against the LiDAR map [8], the recent advancement of Deep Neural Network (DNN)-based methods for multimodal representation learning [9, 10] has enabled the matching to be performed directly in the feature space.
@@ IV-G Monocular Localization
| Camera Dependent | CMRNet [18] | 45 | 1.35 | 60 | 1.35 | - | - | - | - |
@@
| Camera Agnostic | CMRNet++ [24] | 44 | 1.13 | 61 | 1.13 | 77 | 1.01 | 84 | 2.23 |
@@
where , , , and are the ground truth translation, predicted translation, ground truth quaternion, and predicted quaternion, respectively, and are the multiplicative and inverse operations for quaternions. The results reported in Table II show that our proposed CMRNext outperforms all existing methods on all datasets, except for the translation error on KITTI, where I2P-Net achieved a slightly smaller error. However, the rotation error of CMRNext is significantly smaller than that of I2P-Net, Moreover, it is important to note that I2P-Net is trained on KITTI only and therefore more likely to overfit on this dataset. Additionally, CMRNext employs an iterative refinement approach, thus the localization error can be further decreased, surpassing the performance of I2P-Net already at the second iteration, as discussed in Section IV-G1.
@@
| KITTI | Argoverse | Pandaset | ||||
| Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | |
| PWCNet [37] | 44.02 | 1.13 | 60.94 | 1.13 | 77.68 | 1.01 |
| MaskFlowNet-S [61] | 23.18 | 0.57 | 28.86 | 0.38 | 41.28 | 0.44 |
| GMFlow [47] | 27.06 | 0.58 | 87.65 | 1.45 | 61.38 | 0.74 |
| GMFlow (refine) [47] | 11.22 | 0.30 | 33.04 | 0.44 | 22.19 | 0.24 |
| FlowFormer [62] | 15.25 | 0.36 | 81.17 | 1.02 | 30.37 | 0.28 |
| FlowFormer (tiled) [62] | 10.34 | 0.30 | 16.04 | 0.24 | 17.22 | 0.20 |
| RAFT [46] | 9.55 | 0.29 | 12.72 | 0.20 | 12.43 | 0.15 |
| Positional Encoding | Occlusion Filter | Voxel Size [m] | Context Encoder | KITTI | Argoverse | Pandaset | |||
| Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | ||||
| ✓ | 0.1 | RGB | 12.83 | 0.33 | 23.50 | 0.38 | 24.35 | 0.27 | |
| 11.27 | 0.31 | 28.82 | 0.41 | 22.03 | 0.24 | ||||
| 11.02 | 0.30 | 23.49 | 0.30 | 19.91 | 0.23 | ||||
| =12 | 0.1 | RGB | 12.25 | 0.33 | 38.83 | 0.66 | 37.77 | 0.54 | |
| 11.02 | 0.30 | 23.49 | 0.30 | 19.91 | 0.23 | ||||
| =12 | ✓ | RGB | 33.51 | 0.64 | 81.94 | 0.76 | 41.98 | 0.34 | |
| 20.15 | 0.43 | 68.86 | 0.66 | 43.67 | 0.41 | ||||
| 11.02 | 0.30 | 23.49 | 0.30 | 19.91 | 0.23 | ||||
| =12 | ✓ | 0.1 | 11.02 | 0.30 | 23.49 | 0.30 | 19.91 | 0.23 | |
| 9.55 | 0.29 | 12.72 | 0.20 | 12.43 | 0.15 | ||||
@@ IV-I LiDAR-Camera Calibration
| KITTI - left | KITTI - right | |||
| Transl.[cm] | Rot.[°] | Transl.[cm] | Rot.[°] | |
| CMRNet [18] | 1.57 | 0.10 | 52.92 | 1.49 |
| LCCNet [36] | 1.01 | 0.12 | 52.51 | 1.47 |
| CMRNext (ours) | 1.89 | 0.08 | 7.07 | 0.23 |
The left camera was included in the training set of all methods, while the right camera was never seen during training.
@@ IV-J Extrinsic Calibration — Temporal Aggregation
| Freiburg - Car | Freiburg - Quadruped | Freiburg - UAV | ||||
| Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | |
| CMRNext (Single frame) | 12.20 | 0.90 | 23.95 | 1.36 | 12.47 | 0.97 |
| CMRNext (Temporal Aggregation) | 1.89 | 0.19 | 9.70 | 0.21 | 1.84 | 0.23 |

### v3

@@ I Introduction
Localization is an essential task for any autonomous robot, as it is a precursor for subsequent safety-critical tasks such as path planning and navigation. The accuracy of Global Navigation Satellite Systems is not sufficient for most of these tasks, and therefore sensors such as Light Detection And Ranging (LiDAR) and cameras are employed to estimate the pose of the robot. While LiDAR-based localization methods [1, 2] achieve the best performance in terms of accuracy and robustness, the comparatively high price of LiDAR sensors hinders their widespread adoption. On the other hand, cameras are inexpensive and already widely available in consumer vehicles for Advanced Driver Assistance Systems (ADAS) applications. However, the performance of camera-based localization methods [3, 4, 5, 6] is still not comparable to LiDAR-based localization techniques, especially when the lighting and weather conditions during deployment differ from those recorded while mapping.
Consequently, monocular localization in LiDAR maps has gained considerable attention from both academia and industry in recent years, as it combines the high accuracy and robustness of LiDAR-based localization, while only requiring a camera sensor during deployment. The increasing availability of LiDAR maps from mapping companies such as HERE and TomTom has further contributed to the popularity of this approach. Once a LiDAR map is available, either obtained from a mapping company or generated using recent LiDAR Simultaneouos Localization And Mapping (SLAM) algorithms [7, 2, 8], a vehicle can be localized within this map using a single camera. Considering the low cost of cameras, compared to LiDAR sensors, this approach is particularly appealing for large-scale deployment, where the cost of the final system is a key factor. While traditionally this task has been addressed by first generating a 3D representation from the camera image and then matching it against the LiDAR map [9], the recent advancement of Deep Neural Network (DNN)-based methods for multimodal representation learning [10, 11, 12] has enabled the matching to be performed directly in the feature space.
@@ IV-G Monocular Localization
Camera
Dependent |
CMRNet [20] | 45 | 1.35 | 60 | 1.35 | - | - | - | - |
@@
Camera
Agnostic |
CMRNet++ [26] | 44 | 1.13 | 61 | 1.13 | 77 | 1.01 | 84 | 2.23 |
@@
where , , , and are the ground truth translation, predicted translation, ground truth quaternion, and predicted quaternion, respectively, and -1 are the multiplicative and inverse operations for quaternions. The results reported in Table II show that our proposed CMRNext outperforms all existing methods on all datasets, except for the translation error on KITTI, where I2P-Net achieved a slightly smaller error. However, the rotation error of CMRNext is significantly smaller than that of I2P-Net, Moreover, it is important to note that I2P-Net is trained on KITTI only and therefore more likely to overfit on this dataset. Additionally, CMRNext employs an iterative refinement approach, thus the localization error can be further decreased, surpassing the performance of I2P-Net already at the second iteration, as discussed in Section IV-G1.
@@
| KITTI | Argoverse | Pandaset | |||||
| Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | ||
| PWCNet [39] | 44.02 | 1.13 | 60.94 | 1.13 | 77.68 | 1.01 | |
| MaskFlowNet-S [63] | 23.18 | 0.57 | 28.86 | 0.38 | 41.28 | 0.44 | |
| GMFlow [49] | 27.06 | 0.58 | 87.65 | 1.45 | 61.38 | 0.74 | |
| GMFlow (refine) [49] | 11.22 | 0.30 | 33.04 | 0.44 | 22.19 | 0.24 | |
| FlowFormer [64] | 15.25 | 0.36 | 81.17 | 1.02 | 30.37 | 0.28 | |
| FlowFormer (tiled) [64] | 10.34 | 0.30 | 16.04 | 0.24 | 17.22 | 0.20 | |
| RAFT [48] | 9.55 | 0.29 | 12.72 | 0.20 | 12.43 | 0.15 | |
| Positional Encoding | Occlusion Filter | Voxel Size [m] | Context Encoder | KITTI | Argoverse | Pandaset | ||||
| Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | |||||
| ✓ | 0.1 | RGB | 12.83 | 0.33 | 23.50 | 0.38 | 24.35 | 0.27 | ||
| 11.27 | 0.31 | 28.82 | 0.41 | 22.03 | 0.24 | |||||
| 11.02 | 0.30 | 23.49 | 0.30 | 19.91 | 0.23 | |||||
| =12 | 0.1 | RGB | 12.25 | 0.33 | 38.83 | 0.66 | 37.77 | 0.54 | ||
| 11.02 | 0.30 | 23.49 | 0.30 | 19.91 | 0.23 | |||||
| =12 | ✓ | RGB | 33.51 | 0.64 | 81.94 | 0.76 | 41.98 | 0.34 | ||
| 20.15 | 0.43 | 68.86 | 0.66 | 43.67 | 0.41 | |||||
| 11.02 | 0.30 | 23.49 | 0.30 | 19.91 | 0.23 | |||||
| =12 | ✓ | 0.1 | 11.02 | 0.30 | 23.49 | 0.30 | 19.91 | 0.23 | ||
| 9.55 | 0.29 | 12.72 | 0.20 | 12.43 | 0.15 | |||||
@@ IV-J Extrinsic Calibration — Temporal Aggregation
| Freiburg - Car | Freiburg - Quadruped | Freiburg - UAV | |||||
| Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | ||
| CMRNext (Single frame) | 12.20 | 0.90 | 23.95 | 1.36 | 12.47 | 0.97 | |
| CMRNext (Temporal Aggregation) | 1.89 | 0.19 | 9.70 | 0.21 | 1.84 | 0.23 | |